OpenStudy (anonymous):

The mass of a solid cube is 905 g, and each edge has a length of 5.05 cm. Find the density of the cube. Answer in units of kg/m 3

OpenStudy (bibby):

\(\huge d=\frac{m}{v}\) m=? v=?

OpenStudy (anonymous):

i have tried using that but the only thing i need to know is how do i convert 5.05cm into m^3?

OpenStudy (bibby):

you can't. you can convert cm^3 to m^3 though

OpenStudy (anonymous):

so am i just supposed to cube cm and then convert it to meters^3?

OpenStudy (bibby):

well first, you have to get the volume of a cube with side length 5.05cm, what would that be?

OpenStudy (anonymous):

127.79?

OpenStudy (bibby):

I got 128.787625

OpenStudy (anonymous):

hmm i must have messed up then. so then what would i do with that number?

OpenStudy (bibby):

we use dimensional analysis to get it into meters cubed. \(\large \frac{905g}{128.787625cm^3}=\frac{905g}{128.787625\cancel{cm}*\cancel{cm}*\cancel{cm}} *\frac{100\cancel{cm}*100\cancel{cm}*100\cancel{cm}}{1m*1m*1m}\)

OpenStudy (anonymous):

and what would happen next?

OpenStudy (bibby):

well what we have no is in terms of g/m^3 but we want kg/m^3

OpenStudy (anonymous):

so convert g into kg?

OpenStudy (bibby):

yeah.

OpenStudy (anonymous):

ok give me a sec.

OpenStudy (anonymous):

ok so i ended up with .905 kg/ 128.787625 m^3 and i just do the division?

OpenStudy (bibby):

not exactly. Recall that we have to multiply by 100 3 times to get it into m^3

OpenStudy (anonymous):

i mean yeah i remember

OpenStudy (bibby):

\(\large \frac{905g}{128.787625cm^3}=\frac{905g}{128.787625\cancel{cm}*\cancel{cm}*\cancel{cm}} \) \(\large \frac{905\cancel{g}}{128.787625\cancel{cm}*\cancel{cm}*\cancel{cm}} *\frac{100\cancel{cm}*100\cancel{cm}*100\cancel{cm}}{1m*1m*1m}*\frac{1kg}{1000\cancel{g}}\) is the full work

OpenStudy (anonymous):

so then i would do 905000000 kg/128787.625m^3 ?

OpenStudy (bibby):

well let's simplify it using algebra. we have 1000 in the denominator so we can cancel one 100 and a 10. \[\large \frac{905\cancel{g}}{128.787625\cancel{cm}*\cancel{cm}*\cancel{cm}} *\frac{100\cancel{cm}*10\cancel{100}\cancel{cm}*\cancel{100}\cancel{cm}}{1m*1m*1m}*\frac{1kg}{\cancel{1000}\cancel{g}}\]

OpenStudy (bibby):

poopq

OpenStudy (bibby):

\[\frac{905\cancel{g}}{128.787625\cancel{cm}*\cancel{cm}*\cancel{cm}} *\frac{100\cancel{cm}*10\cancel{100}\cancel{cm}*\cancel{100}\cancel{cm}}{1m*1m*1m}*\frac{1kg}{\cancel{1000}\cancel{g}}\]

OpenStudy (anonymous):

so 905000 kg/ 128.787625 m^3?

OpenStudy (bibby):

which is less ugly if we write \(\huge \frac{905,000kg}{128.787625m^3}\) yeah

OpenStudy (anonymous):

ohhhhh i got 7027.072671. would any more math have to be done or is that it?

OpenStudy (bibby):

http://www.wolframalpha.com/input/?i=7.0270726g%2Fcm%5E3+to+kg%2Fm%5E3 no. wolfram agrees with us

OpenStudy (anonymous):

yessss thank you so much i have been stuck on this problem all day for ap physics thanks:)!

OpenStudy (bibby):

wat this feels chem-y oh well, np np